OpenStudy (anonymous):

Hi, could someone tell me how they got an answer of: x/12 for the question below: A boat costs x dollars, and this cost is to be shared equally by a group of people. In terms of x, how many dollars less will each person contribute if there are 4 people in the group instead of 3.

OpenStudy (anonymous):

at first, there are 3 people so, each pays x/3 dollars after that, there are 4 people so, each pays x/4 dollars the difference between x/3 and x/4 is the amount they pay less =x/3-x/4=x/12

OpenStudy (anonymous):

x/3 - x/4 =4x/12 -3x/12 =(4x-3x)/12 =x/12